Suppose that 45% and 40% of Singaporeans approve and diapprove
the idea of building a casino respectively.If three Singaporeans
are interviewed at random,find the prob that
a)all of them will either approve or disapprove of the idea
b)at least two of them will not approve of the idea
c) at most one of them will approve of the idea.
Let A=approve.DA=disapprove,N=neither approve nor disapprove
p(A)= 0.45,P(DA)=0.40,P(N)=1-0.45-0.40=0.15
a) P(all approve or disapprove)=P(AAA)+P(DA.DA.DA)
b) P(at least 2 of them will not
approve)=P(DA.DA.A)+P(DA.DA.DA)+P(DA.DA.N)=0.40×0.40×0.45+0.40×0.40×0.40+0.40×0.40×0.15
=0.16 but correct answer is 44/125
c)P(at most one of them will
approve)=P(A.DA.DA)+P(A.DA.N)+P(A.N.DA)+P(DA.DA.DA)+P(N.N.N)=0.45×0.40×0.40+0.45×0.40×0.15+0.45×0.15×0.40+0.40×0.40×0.40+0.15×0.15×0.15=0.193375
but correct answer is 2299/4000.
Can someone plse advise..thanks in advance
stats is nightmare for me